OpenStudy (anonymous):

What is the probability of getting either a black card or an ace in one draw from an ordinary deck of 52cards?

OpenStudy (anonymous):

P(black card or ace) = P(black card) + P(ace) - P(black and ace) = 26/52 + 4/52 - 2/52 = 7/13